# GenToC: Leveraging Partially-Labeled Data for Product Attribute-Value Identification

### v1

@@ Abstract.
In the e-commerce domain, the accurate extraction of attribute-value pairs from product listings (e.g., Brand: Apple) is crucial for enhancing search and recommendation systems. The automation of this extraction process is challenging due to the vast diversity of product categories and their respective attributes, compounded by the lack of extensive, accurately annotated training datasets and the demand for low latency to meet the real-time needs of e-commerce platforms. To address these challenges, we introduce GenToC, a novel two-stage model for extracting attribute-value pairs from product titles. GenToC is designed to train with partially-labeled data, leveraging incomplete attribute-value pairs and obviating the need for a fully annotated dataset. The first stage uses a marker-augmented generative model to propose potential attributes, followed by a token classification model that determines the associated values for each attribute. GenToC outperforms existing state-of-the-art models, exhibiting upto 56.3% increase in the number of accurate extractions.
Moreover, we introduce a bootstrapping method that enables GenToC to progressively refine and expand its training dataset. This enhancement substantially improves the quality of data available for training other neural network models that are typically faster but are inherently less capable than GenToC in terms of their capacity to handle partially-labeled data. By supplying an enriched dataset for training, GenToC significantly advances the performance of these alternative models, making them more suitable for real-time deployment. Our results highlight the unique capability of GenToC to learn from a limited set of labeled data and to contribute to the training of more efficient models, marking a significant leap forward in the automated extraction of attribute-value pairs from product titles. GenToC has been successfully integrated into India’s largest B2B e-commerce platform, IndiaMART.com, achieving a significant increase of 21.1% in the number of correctly identified attribute-value pairs over the existing deployed system while achieving a high precision of 89.5%.
@@ 1. Introduction
The rapid expansion of E-Commerce has led to a significant increase in the variety and complexity of products available online. Every product usually comes with a variety of attributes like Brand, Color, Model and so on, each having specific values, for example, Samsung, Phantom Gray, Galaxy S21, etc. (as illustrated in Table 1). These attributes and values help consumers locate and select their desired products. Nonetheless, the attribute-value pairs provided in product listings by sellers are often incomplete and can be enhanced using the information contained within the product title. This problem of automatic attribute-value identification for E-Commerce products is a well-established topic in literature (Bing et al., 2012; Putthividhya and Hu, 2011; Probst et al., 2007; Shinzato and Sekine, 2013; Roy et al., 2021; Wang et al., 2020; Xu et al., 2019; Zheng et al., 2018). It has been investigated in various contexts that sometimes include additional metadata such as product descriptions (Yang et al., 2021; Shinzato et al., 2023), knowledge graphs (Ricatte and Crisostomi, 2023) or images (Zhu et al., 2020; Khandelwal et al., 2023). In this study, we focus on automated extraction from product titles.
Product Name: Boat Rockerz 255 Pro | |
@@
The predominant approach for acquiring training datasets for attribute-value extraction tasks typically relies on leveraging product listings and associated attribute-value pairs as furnished by vendors on E-Commerce platforms. This method is directly dependent on the comprehensiveness and accuracy of the data that vendors provide. However, it is common practice for vendors to list these details in an incomplete or inconsistent manner. Such gaps and irregularities in the data present a substantial obstacle, as the development of effective deep learning models for attribute-value extraction is contingent upon the availability of large-scale, high-quality training data.
The complexity of the task is further amplified by the multitude of attributes involved, potentially running into tens of thousands of attributes (24.7K in our case). At this scale, some attributes end up being redundant, noisy or rarely used which leads to an extremely long-tailed distribution (Shinzato et al., 2023; Xu et al., 2019). This redundancy is evident when attributes that fundamentally represent the same product feature – Color and Shade or Size and Dimensions – are utilized interchangeably. Additionally, attribute representation may vary, with instances such as Model No. and Model Number being treated as distinct attributes despite their equivalence. This adds complexity to the process of training and evaluating neural models for extracting attribute-value pairs, as these overlapping attributes may differ across various products and categories. Finally, attribute-value extraction systems are also often used on user query traffic to enable capturing user search intentions effectively (Ricatte and Crisostomi, 2023). Therefore, developing systems that deliver responses in real-time is crucial.
@@
To address the three challenges outlined, we propose a novel two-stage model, GenToC, which decomposes the task into two distinct phases: initially, it identifies all attributes within a product name, and subsequently, it extracts the corresponding values for each attribute. Specifically, GenToC employs a Generative Seq2Seq model in the initial stage and a Token Classification model in the subsequent stage to accurately extract attribute-value pairs from product names. The first-stage generative model is designed to output all relevant attributes concatenated with each other using a delimiter. The second-stage token classification model takes as input, each attribute name and the product name and classifies the words in the product name that denote the attribute value.
To foster the model’s learning capacity using partially-labeled data, we incorporate special markers during the training phase of the first-stage Generative model, which generates all the attributes potentially present in the product name. These markers are used to highlight the words in the product name that are present as attribute values in training data. For instance, in Table 1, markers will be added to the words Boat, Rockerz, 255, Pro, Raging, Red (at token level) because these are identified as attribute values in the training data. The markers serve as cues, directing the model to assimilate information from these specified words into the attribute generation process while still taking into account the full context of the product name. Through the strategic use of these markers, the model is trained to form associations between attributes and the highlighted words, while also understanding that words without markers do not have any attributes associated with them at training.
Through this training methodology, the model becomes adept at generalization during inference, where we apply a marker to every token in the input. Marking each token prompts the model to treat every word as a potential source of attribute information. This practice significantly refines the model’s predictive strategy, as it becomes more attentive to the potential that any word in a product name might be linked to an attribute. Consequently, the model is not limited to merely mimicking the patterns observed in training but is also capable of anticipating attributes for every word in the product name. Such a strategy is particularly effective when the model is trained on data that is only partially labeled, enabling it to reliably extract attribute information.
The second-stage token classification model accepts both the attribute name and the product name as input, concatenating them with a special delimiter (¡sep¿). It is tasked with classifying each word within the product name to ascertain whether it signifies a value for the given input attribute. In the training phase, we use the attributes that are already known, whereas during inference, we rely on the attributes predicted by the first-stage model. By using the attribute name directly, we enable synonymous attributes to be embedded closely in the vector space. This method effectively addresses the second challenge of the model learning independently from each of the closely associated attributes. So, the training examples for Model No. will influence those for Model Number owing to the similarity in both their semantic representation.
@@
Due to the cascading design of GenToC, errors from the first-stage model can potentially escalate, adversely affecting the second-stage model’s performance. To fortify the second-stage model against such compounded inaccuracies, we implement a ‘Value Pruning’ method. This technique equips the value-extraction model with the ability to generate null outputs for any incorrect attributes that might be output by the attribute-extraction model. Compared to existing state-of-the-art NER models (Zheng et al., 2018) and Generative models (Shinzato et al., 2023) used for attribute-value extraction, GenToC achieves an increase of 16.2% and 18.2% in F1, respectively.
Motivated by the objective of creating a faster model, we employ GenToC to refine the training dataset by predicting the attribute-value pairs from product names within the training data. This bootstrapping process yields a more exhaustive training set. Utilizing this augmented dataset to train the faster NER model results in a rise of 20.2% in F1, compared to training on the original dataset, while giving a real-time speed of 8.8 ms per query during inference.
@@
Introduction of GenToC, a novel two-stage architecture integrating a Generative Seq2Seq with a Token Classification model, leading to an F1 score improvement of 15.5% to 19.7% over existing attribute-value extraction techniques.
Implementation of markers in the initial generative model stage to effectively handle the challenge of incomplete attribute-value pair data, a common issue in E-Commerce product listings used to create training data.
Usage of attribute names in second-stage token classification model that allows to mitigate issues of redundancy and noise across thousands of attributes.
Adoption of bootstrapping to refine training datasets, which leads to strong performance of real-time systems.
@@ 2. Related Work
Attribute value extraction (Nadeau and Sekine, 2007; Bing et al., 2012; Putthividhya and Hu, 2011; Probst et al., 2007; Shinzato and Sekine, 2013; Roy et al., 2021; Wang et al., 2020; Xu et al., 2019; Zheng et al., 2018) has been a significant topic of study in the realm of E-Commerce, with a wealth of research targeting the extraction of product details from various modalities, including purely text-based methods (Wang et al., 2020; Xu et al., 2019) as well as those that incorporate images (Zhu et al., 2020; Khandelwal et al., 2023; Wang et al., 2023). A common limitation of these methods is their assumption of high-quality training data, which fails in many real-world scenarios where the training data is created using distant supervision and hence potentially incomplete (Shrimal et al., 2022; Xu et al., 2019). Some publicly available datasets like MAVE (Yang et al., 2021) offer high-quality data (achieving over 98% F1 score), but our work specifically addresses the more challenging case of partially-labeled settings, and therefore, we do not compare with them.
@@
Few works, such as Zhang et al. (2021); Zheng et al. (2018) focus on improving the training data quality in attribute-value extraction task but are severely limited to operating on a small number of attributes only. For instance, Zhang et al. (2021) rely on a subset of strongly-labeled data to train a teacher network which in turn creates training data for a student network. The requirement of having a strongly-labeled subset limits them to operate on 13 attributes. Zheng et al. (2018) also relies on a small set of labeled instances to be used in an active learning setting to collect good examples for manual annotation. They apply the technique to only one attribute per dataset. In contrast, GenToC introduces a fundamentally new model design that can handle incomplete training data without requiring any completely labeled subsets. This allows the model to scale to tens of thousands of attributes.
@@
In terms of modelling, some of the earliest works on attribute value extraction primarily employed rule-based extraction methods. They utilized a specialized seed dictionary or vocabulary to identify key phrases and attributes (Ghani et al., 2006; Wong et al., 2009; Gopalakrishnan et al., 2012). Moving beyond rule-based systems, several neural models have also been proposed for this task in the recent past (Roy et al., 2021). Architectures for this task can be broadly divided into two categories – token classification (Yang et al., 2021; Xu et al., 2019; Shrimal et al., 2022) or generative (Roy et al., 2022; Shinzato et al., 2023) models. Token classification models typically identify the spans in the input text corresponding to an attribute. On the other hand, generative models utilize Seq2Seq models to produce relevant attribute-value pairs from a specified input. Our GenToC model makes clever use of both the types of architecture, using a generative model for attribute extraction and a token-classification model for value extraction. We compare extensively with both types of architectures and show significant gains achieved by GenToC.
@@
Token classification systems that rely on NER are often limited to the attributes seen in training time (Zhang et al., 2021). Other token-classification systems utilize a BERT+LSTM model to embed each attribute separately (Xu et al., 2019). The embedded representation attends over the product name and identifies the corresponding value. This allows the inclusion of new attributes at the inference stage. However, this approach encounters scalability issues as the model must check for values using every possible attribute. Some generative models (Khandelwal et al., 2023) are limited in their operation to a relatively small set of attributes (in their case, 38). This is because they use a seq2seq model pass for every attribute to identify the relevant values. Other generative models (Shinzato et al., 2023), are applicable to larger attribute sets as they generate the attribute name as well. Similarly, GenToC can handle large attribute sets due to a dedicated module that generates all the relevant attributes.
@@ 3. Problem Statement
Given a product name with words, and the set of all potential attributes, , the objective of attribute-value extraction is to list all possible attribute-value pairings, which can be denoted as , where and is a subset of words in . In our setup, we assume that each word in is linked to a maximum of one attribute.
4. Models
In this section, we first describe how existing State-of-the-art models use Generative models (Gen-AVE) and Token Classification (ToC-AVE) for attribute-value extraction. We then proceed with introducing GenToC that combines both generative and token classification in the two stages of the model. The three models are visually represented in Figure 1, and comparative analysis is presented in Table 2.
4.1. Generative Attribute-Value Extraction
Following the architecture proposed in Shinzato et al. (2023), we design the Gen-AVE model to employ a Seq2Seq model that inputs the product name and yields a concatenated string incorporating the respective attribute-value pairs. For instance, the encoder takes an input , and the decoder generates the output string ‘:,:,, :’. Its challenges are slow inference speed and limited learning ability from partially labeled data.
4.2. Token Classification Attribute-Value Extraction
@@
Following NER models used for the task (Rezk et al., 2019; Shinzato et al., 2022), ToC-AVE is designed as an encoder-only model that operates by classifying tokens in the product name, where each token in the query is assigned a label corresponding to the relevant attribute. As a NER model, it treats every attribute as a unique label and is limited to the attributes it has seen during training. However, its major strength lies in its speed due to the simple encoder-only architecture. Although it is not capable of effectively learning from partially labeled data, we find that training it with better quality data generated by GenToC, results in a fast and powerful attribute-value extraction system.
4.3. GenToC: Generative Attribute Extraction + Token Classification Value Extraction

### v2

# A Framework for Leveraging Partially-Labeled Data for Product Attribute-Value Identification

@@ Abstract.
In the e-commerce domain, the accurate extraction of attribute-value pairs (e.g., Brand: Apple) from product titles and user search queries is crucial for enhancing search and recommendation systems. A major challenge with neural models for this task is the lack of high-quality training data, as the annotations for attribute-value pairs in the available datasets are often incomplete. To address this, we introduce GenToC, a model designed for training directly with partially-labeled data, eliminating the necessity for a fully annotated dataset. GenToC employs a marker-augmented generative model to identify potential attributes, followed by a token classification model that determines the associated values for each attribute. GenToC outperforms existing state-of-the-art models, exhibiting upto 56.3% increase in the number of accurate extractions. Furthermore, we utilize GenToC to regenerate the training dataset to expand attribute-value annotations. This bootstrapping substantially improves the data quality for training other standard NER models, which are typically faster but less capable in handling partially-labeled data, enabling them to achieve comparable performance to GenToC. Our results demonstrate GenToC’s unique ability to learn from a limited set of partially-labeled data and improve the training of more efficient models, advancing the automated extraction of attribute-value pairs. Finally, our model has been successfully integrated into IndiaMART, India’s largest B2B e-commerce platform, achieving a significant increase of 20.2% in the number of correctly identified attribute-value pairs over the existing deployed system while achieving a high precision of 89.5%.
@@ 1. Introduction
The rapid expansion of e-commerce has led to a significant increase in the variety and complexity of products available online. Each product typically includes a set of attributes such as Brand, Model Name, Color, with distinct values like Boat, Rockerz 255 Pro, Raging Red (as demonstrated in Table 1). These attributes and values help consumers locate and select their desired products. Automatic attribute-value identification is a well-studied problem in the e-commerce literature (Bing et al., 2012; Putthividhya and Hu, 2011; Probst et al., 2007; Shinzato and Sekine, 2013; Roy et al., 2021; Wang et al., 2020; Xu et al., 2019; Zheng et al., 2018). It has been investigated in various contexts that involve additional metadata, such as product descriptions (Yang et al., 2021; Shinzato et al., 2023), knowledge graphs (Ricatte and Crisostomi, 2023) or images (Zhu et al., 2020; Khandelwal et al., 2023). In this study, we focus on automated extraction using only textual information (Xu et al., 2019; Zheng et al., 2018), such as product titles and user search queries.
To illustrate the practical utility of attribute-value extraction systems, we examine their usage in IndiaMART111https://www.indiamart.com, India’s largest B2B e-commerce platform, where our system is currently deployed. In this platform, our attribute-value extraction system serves a dual purpose, enhancing both product listings and user searches. The attribute-value pairs provided in product listings by sellers are often incomplete and can be enhanced using extractions from the product title. In addition, the system is used to extract attribute-value pairs from user search queries, and these extracted pairs are matched with attribute-value pairs from the retrieved product listings to highlight the matching ones. This dynamic feature highlighting of product attribute-value pairs based on the search query enables users to easily find the products that meet their requirements.
Product Title: Boat Rockerz 255 Pro | |
@@
The predominant approach for acquiring training datasets for attribute-value extraction tasks typically relies on leveraging product listings and associated attribute-value pairs as furnished by vendors on e-commerce platforms. This method is directly dependent on the comprehensiveness and accuracy of the data that vendors provide. However, it is common practice for vendors to list these details in an incomplete or inconsistent manner. Such gaps and irregularities in the data present a substantial obstacle, as the development of effective deep learning models for attribute-value extraction is contingent upon the availability of large-scale, high-quality training data. Given that attribute-value extraction systems are also employed on user query traffic to accurately capture user search intentions (Ricatte and Crisostomi, 2023), it is essential to develop light-weight systems that provide real-time responses while managing the substantial workloads prevalent in e-commerce platforms.
To address these dual challenges of incomplete training data and the need for real-time responses, we propose a novel framework. First, we use a specialized model designed to learn from the incomplete data. This model is then used to regenerate the training data, thereby enabling the training of a faster model that can deliver real-time responses. This overall framework is illustrated in Figure 1.
@@
To address the challenges of incomplete training data, we introduce GenToC, a novel model that uses a specialized marker learning strategy. GenToC is a two-stage neural model that decomposes the task into two distinct phases. First, it identifies all attributes within the input text, which can be either a product title or a user search query. Then, it extracts the corresponding values for each attribute. Specifically, GenToC employs a Generative Seq2Seq model in the initial stage and a Token Classification model in the subsequent stage to accurately extract attribute-value pairs from the input text. The first-stage generative model is designed to output all relevant attributes concatenated with each other using a delimiter. The second-stage token classification model receives each attribute name along with the original input query and determines for each word in the text whether it belongs to the value of the input attribute. For example, for the input text in Table 1, the first stage model can identify attributes such as Brand, Model Name, and Color. Then, the second stage model assigns Boat, Rockerz 255 Pro and Raging Red as the values for the corresponding attributes.
To foster the model’s learning capacity using partially-labeled data, we incorporate special markers during the training phase of the first-stage generative model. These markers are used to highlight the words in the input query that are present as attribute values in the training data. For example, in Table 1, markers will be added to the words Boat, Rockerz, 255, Pro, Raging, Red (at token level) because these are identified as attribute values in the training data. Through the strategic use of these markers, the model learns to form associations between attributes and highlighted words, that every highlighted word is a potential source of attribute information. This helps the model during inference, where we apply a marker to every token in the input. Applying a marker to each token serves as a signal to the model to regard every word as a possible source of attribute information, since similar correlations were observed during training. Thus, the model tends to produce a wider range of attributes and is not limited to mimicking the incomplete patterns observed in training. Existing conventional models lack this generalization capability, making it difficult to utilize partially labeled data effectively.
@@
The second-stage token classification model accepts both the attribute name and the original query as input. It is tasked with classifying each word within the query to ascertain whether that signifies a value for the given input attribute. In the training phase, we use the attributes that are already known, whereas during inference, we rely on the attributes predicted by the first-stage model. Due to the cascading design of GenToC, errors from the first-stage model can potentially escalate, adversely affecting the second-stage model’s performance. To fortify the second-stage model against such compounded inaccuracies, we implement a ‘Value Pruning’ method. This technique equips the value-extraction model with the ability to generate null outputs for any incorrect attributes that might be output by the first-stage attribute-extraction model.
Compared to existing state-of-the-art NER models (Zheng et al., 2018) and Generative models (Shinzato et al., 2023) used for attribute-value extraction, GenToC achieves an increase of 16.2% and 18.2% in F1-score, respectively, on the test set. Despite these substantial performance gains, its speed of 90 ms per query presents a limitation for deployment.
Motivated by the objective of creating a faster model, we use GenToC to predict attribute-value pairs for each query in the training data and regenerate the dataset. This bootstrapping process improves attribute-value tagging in the training dataset. By training an NER model using this augmented dataset, we achieve an increase of 16.8% in F1-score compared to using the original incomplete dataset. It is also now comparable to GenToC in terms of F1-score, while maintaining an inference speed of under 9 ms per query. The GenToC-bootstrapped NER system is currently deployed on IndiaMART, replacing a rule-based system and has already served over 200M requests. Compared to the previous system, our new system has improved the recall of attribute-value extraction by over 20% on product titles and has resulted in a 9% increase in user search queries that activate dynamic feature highlighting.
@@
Introducing GenToC, a novel two-stage architecture featuring a Generative Seq2Seq model for attribute extraction followed by a Token Classification model for value mapping, scalable to tens of thousands of attributes.
Utilizing markers-based learning in GenToC’s first stage to handle incomplete attribute-value tagging in training data.
Surpassing existing methods with a 16.2% F1-score increase.
Enhancing the training data using bootstrapping, thereby boosting the F1-score of NER model trained with it by 16.8%.
Deploying the system in a leading B2B platform, where it has served over 200M requests in production so far.
@@ 2. Related Work
Attribute-value extraction (Nadeau and Sekine, 2007; Bing et al., 2012; Putthividhya and Hu, 2011; Probst et al., 2007; Shinzato and Sekine, 2013; Roy et al., 2021; Wang et al., 2020; Xu et al., 2019; Zheng et al., 2018) has been a significant topic of study in the realm of e-commerce, with a wealth of research targeting the extraction of product details from various modalities, including purely text-based methods (Wang et al., 2020; Xu et al., 2019) as well as those that incorporate images (Zhu et al., 2020; Khandelwal et al., 2023; Wang et al., 2023). A common limitation of these methods is their assumption of high-quality training data, which fails in many real-world scenarios where the training data is created using distant supervision and hence potentially incomplete (Shrimal et al., 2022; Xu et al., 2019). Some publicly available datasets like MAVE (Yang et al., 2021) offer high-quality data (achieving over 98% F1-score), but our work specifically addresses the more challenging case of partially-labeled settings, where existing models struggle.
@@
Few works, such as Zhang et al. (2021); Zheng et al. (2018) focus on improving the training data quality in attribute-value extraction tasks but are severely limited to operating on a small number of attributes only. For instance, Zhang et al. (2021) rely on a subset of strongly-labeled data to train a teacher network which in turn creates training data for a student network. The requirement of having a strongly-labeled subset limits them to operate on 13 attributes. Zheng et al. (2018) also relies on a small set of labeled instances to be used in an active learning setting to collect good examples for manual annotation. They apply the technique to only one attribute per dataset. In contrast, GenToC introduces a fundamentally new model design that can handle incomplete training data without requiring any completely labeled subsets. This allows the model to scale to tens of thousands of attributes.
@@
In terms of modelling, some of the earliest works on attribute-value extraction primarily employed rule-based extraction methods. They utilized a specialized seed dictionary or vocabulary to identify key phrases and attributes (Ghani et al., 2006; Wong et al., 2009; Gopalakrishnan et al., 2012). Moving beyond rule-based systems, several neural models have also been proposed for this task in the recent past (Roy et al., 2021). Neural architectures for this task can be broadly divided into two categories – token classification (Yang et al., 2021; Xu et al., 2019; Shrimal et al., 2022) or generative (Roy et al., 2022; Shinzato et al., 2023) models. Token classification models utilize NER models to identify the spans in the input text corresponding to an attribute. On the other hand, generative models utilize Seq2Seq models to produce relevant attribute-value pairs from a specified input. Our GenToC model makes clever use of both types of architectures, using a generative model for attribute extraction and a token-classification model for value extraction. We compare extensively with both types of architectures and show significant gains achieved by GenToC.
@@
Token classification systems that rely on NER use atomic embeddings for encoding attributes, which makes it challenging to handle long-tail and complex attributes with less training data. Other token-classification systems utilize a BERT+LSTM model to embed each attribute separately (Xu et al., 2019). The embedded representation attends over the input query and identifies the corresponding value. However, this approach encounters scalability issues as the model must check for values using every possible attribute. Some generative models (Khandelwal et al., 2023) are limited in their operation to a relatively small set of attributes (in their case, 38). This is because they use a Seq2Seq model pass for every attribute to identify the relevant values. Other generative models (Shinzato et al., 2023) are applicable to larger attribute sets as they generate the attribute name as well. Similarly, GenToC can handle large attribute sets due to a dedicated module that generates all the relevant attributes. Like other generative models, GenToC employs compositional encoding for attributes, which allows it to capture the semantic meaning of attribute names by considering the words they contain. So it is better at handling long-tail and complex attributes compared to NER models. But the generative nature of the model comes with a higher response time. We summarize the characteristics of the different models in Table 2.
Model | Partially labeled | Long-tail attributes | Response time |
NER | ✗ | ✗ | fast |
Seq2Seq | ✗ | ✓ | slow |
GenToC | ✓ | ✓ | slow |
@@ 3. Problem Statement
Given a query with words, where , and a set of all potential attributes , the objective of attribute-value extraction is to identify all possible attribute-value pairs, denoted as , where and is a subset of words in . In our setup, the input may be a product title or user search query, with each word in linked to at most one attribute.
4. Background
In this section, we describe in detail two popular architectures that have been commonly used for the task of attribute-value extraction. We illustrate their working in Figure 2 (a) and (b).
4.1. NER-AVE
@@
Following NER models used for the task (Rezk et al., 2019; Shinzato et al., 2022), NER-AVE is an encoder-only model that operates by classifying tokens in the input query, where each token in the query is assigned a label corresponding to the relevant attribute. For example, boat is assigned the attribute Brand, each of three words rockerz, 255 and pro are assigned the attribute Model Name, and so on. In case no attribute exists for a particular word, a special label, NoAttribute is used to indicate the same. As an NER model, it treats every attribute as a unique label and assigns an atomic embedding. However, its major strength lies in its speed due to the simple encoder-only architecture. Although it is not capable of effectively learning from partially labeled data, we find that training it with better quality data generated by GenToC, results in a fast and powerful attribute-value extraction system.
4.2. Seq2Seq-AVE
The Seq2Seq-AVE model is based on the model developed in Shinzato et al. (2023), which employs a Seq2Seq model that yields a concatenated string, incorporating the respective attribute-value pairs, for a given input query. For instance, the encoder takes an input , and the decoder generates the output string ‘:, :, , :’. For example, the decoder produces the generation “Brand: boat, Model Name: rockerz 255 pro …”, one token at a time. The attribute and value pairs can then be parsed from this generation. Its limitations are slow inference speed (found to be over 10x slower than NER-AVE) and limited learning ability from partially labeled data.
5. Methods